OpenStudy (anonymous):

Shannon opens a savings account with $50. Starting with the first week, each week she deposits $8 into her account. How many weeks does it take her to save $450? A. 56.25 weeks B. 62.5 weeks C. 52 weeks D. 50 weeks

OpenStudy (ranga):

50 + N * 8 = 450 (where N is the number of weeks to save $450). Solve for N.

OpenStudy (anonymous):

You start with 50, so subtract that. It leaves you with 400. Divide that by 8 and you have your answer.

OpenStudy (sleepyjess):

50 + (x*8) = 450 50-50 + (x*8) = 450-50 x*8 = 400 x*8/8 = 400/8 x =